# Calibration Method of Spacecraft–Inertial Sensor Center-of-Mass Offset for the Taiji Gravitational Wave Detection Mission under Science Mode

### v1

@@ III.1 Controller design
Building upon the control requirements and SC–TM dynamics established in Sections I and II, this work adopts the DFACS architecture proposed by [vidano2024ieee] as the baseline design.
@@
The $H_{\infty}$ control method is a robust synthesis approach rooted in norm-optimal control theory, specifically formulated to ensure closed-loop stability and prescribed performance bounds in the presence of model uncertainties and exogenous disturbances. As illustrated in Figure 2, the generalized plant $P$ represents the controlled dynamics (e.g., SC–TM coupled system), while $K$ denotes the to-be-designed controller. The objective is to synthesize a stabilizing controller $K$ that minimizes the worst-case gain from exogenous inputs $\mathbf{w}$ (e.g., disturbances, sensor noise) to regulated outputs $\mathbf{z}$ (e.g., TM–SC relative motion, attitude error).
This is formalized as the $H_{\infty}$ norm minimization problem [gu2005robust]
@@
Since analytical $H_{\infty}$ solutions are generally intractable [gu2005robust], a suboptimal controller is obtained by minimizing $\eta$ subject to
@@
which reduces to a convex linear matrix inequality (LMI) feasibility problem via the Bounded Real Lemma [LMIboyd1994linear].
The suboptimal $H_{\infty}$ problem (16) can be solved by Mixed Sensitivity Design [ROBUST_CONTROL] . Its inputs are the plant $G$ and weighting functions $W_{1}$, $W_{2}$, $W_{3}$, representing performance, control effort, and robustness requirements, respectively. The synthesis ensures the frequency-dependent bounds
@@
Where the sensitivity and complementary sensitivity functions are defined as $S=(I+GK)^{-1}$ and $T=GK(I+GK)^{-1}$. Enforcing an upper bound on $\|S\|$ improves low-frequency disturbance rejection and reference tracking; bounding $\|T\|$ ensures high-frequency robustness against sensor noise and unmodeled dynamics; and limiting $\|KS\|$ constrains control effort. Since $H_{\infty}$ synthesis via mixsyn requires a state-space realization, the nonlinear DFACS dynamics, i.e, Eq. (4), Eq. (7), and Eq. (11) are linearized about the nominal drag-free trajectory. The resulting linear time-invariant (LTI) model is
@@
The specific mathematical model is similar to that of [vidano2024drag], where $\mathbf{x}=[\mathbf{p}^{\top}\ \mathbf{\dot{p}}^{\top}]^{\top}$ is the state vector comprising relative position $\mathbf{p}$ and velocity $\mathbf{\dot{p}}$. The state variables of position are defined as
@@
where $\mathbf{F}_{E_{i}}\in\mathbb{R}^{2}$ denotes the 2-D electrostatic force on the $i$-th TM in the plane orthogonal to its sensitive axis.
@@
where $\mathbf{B}_{v}^{\dagger}\in\mathbb{R}^{16\times 15}$ denotes the pseudoinverse of the input matrix $\mathbf{B}_{v}\in\mathbb{R}^{15\times 16}$. Substituting (21) into the velocity dynamics $\ddot{\mathbf{p}}=\mathbf{A}_{v}\mathbf{p}+\mathbf{B}_{v}\mathbf{u}$ yields the fully decoupled closed-loop form
@@
To satisfy the drag-free control requirements, MPS is commanded to allocate its thrust in the SRF as follows: the $x$-axis thrust component generates the control force associated with the symmetric mode $x_{m_{1}O_{1}}^{O_{1}}+x_{m_{2}O_{2}}^{O_{2}}$, while the $y$-axis thrust component drives the antisymmetric mode $x_{m_{1}O_{1}}^{O_{1}}-x_{m_{2}O_{2}}^{O_{2}}$:
$\mathbf{F}_{T,x}$ $=B_{v,{1,4}}^{\dagger}\bigl(u_{d,4}+u_{d,10}\bigr),$
$=B_{v,1,10}^{\dagger}\bigl(u_{d,4}+u_{d,10}\bigr).$ (24)
$\mathbf{F}_{T,y}$ $=B_{v,2,4}^{\dagger}\bigl(u_{d,4}-u_{d,10}\bigr),$
$=B_{v,2,10}^{\dagger}\bigl(-u_{d,4}+u_{d,10}\bigr).$ (25)
Clearly, this allocation strategy imposes constraints on Eq. (23), analogous constraints for attitude channels based on the SC and TM inertia tensors (e.g., torque allocation proportional to $\mathbf{J}_{S}^{-1}$, $\mathbf{J}_{m}^{-1}$). This convex quadratic program is solved using active-set algorithm. The resulting control allocator achieves a projection error of $\|\mathbf{I}_{15}-\mathbf{B}_{v}\mathbf{B}_{v}^{\dagger}\|_{2}=2.2\times 10^{-3}$, indicating near-exact command realization.
After completing the decoupling, the design of the $H_{\infty}$ controller continues. Eq. (21) and Eq. (22) enable us to shift from designing a controller for $\mathbf{u}$ to designing a controller for the decoupled SISO system $\mathbf{u}_{d}$. For each degree of freedom in $\mathbf{p}$, only the associated control requirements and noise characteristics pertinent to that specific degree of freedom need to be considered in determining the weighting functions in Eq. (17). The detailed DFACS control requirements are given in Table 1 [vidano2024drag]:
@@
state variable requirement (ASD)
@@
$y_{m_{i}O_{i}}^{O_{i}},\,z_{m_{i}O_{i}}^{O_{i}}$ $5\times 10^{-9}\,\mathrm{m}/\sqrt{\mathrm{Hz}}$
$\varphi_{SC},\,\theta_{SC},\,\psi_{SC}$ $1\times 10^{-8}\,\mathrm{rad}/\sqrt{\mathrm{Hz}}\times\omega_{b}$
$\varphi_{m_{i}O_{i}}$ $2\times 10^{-7}\,\mathrm{rad}/\sqrt{\mathrm{Hz}}\times\omega_{b}$
$\theta_{m_{i}O_{i}},\,\psi_{m_{i}O_{i}}$ $1\times 10^{-8}\,\mathrm{rad}/\sqrt{\mathrm{Hz}}\times\omega_{b}$
@@
Where $F_{u}$ is the transfer function determined by the exogenous disturbance model, $F_{y}$ is the transfer function determined by the measurement noise model, and $B$ is determined by the control requirements of each state variable. In the disturbance models, special attention is given to solar radiation pressure (SRP), with its transfer function given by [virdis2021meteoroid]
@@
The constant $k_{d\mathrm{SC}}$ is derived from the solar radiation pressure constant at the SC’s orbital distance and the effective area of its solar arrays [SRP_Georgevic1971]. Given the similarity in orbital configuration ($\mathbf{r}_{SI}^{I}\approx 1~\text{AU}$) and spacecraft structural layout, we adopt the value used in the LISA mission studies, namely, $k_{d\mathrm{SC}}=6.3513\times 10^{-5}~\text{N},$ which is significantly larger than the magnitude of the transfer function $H_{\mathrm{press}}$ ($\sim 1\times 10^{-10}~\text{N}$). Therefore, including SRP directly in $F_{u}$ would significantly increase the requirements on $W_{1}$ in the low-frequency range, making Eq. (17) difficult to solve. To address this issue, we introduce a feedforward channel that leverages spacecraft orbit determination data to estimate $\hat{\mathbf{r}}_{SI}^{I}$. Based on Taiji’s exceptional orbital stability and its periodic ground-based orbit determination architecture, a conservative 50-km $1\sigma$ position estimation error is assumed for $\hat{\mathbf{r}}_{SI}^{I}$ [li2021orbit]. This allows for the estimation of the ultra-low-frequency components of $\mathbf{F}_{SRP}$ and $\mathbf{D}_{SRP}$, denoted as $\hat{\mathbf{F}}_{SRP,C}$ and $\hat{\mathbf{D}}_{SRP,C}$
@@
Figure 3 shows the mixed-sensitivity $H_{\infty}$ controller design results. Critically, both the sensitivity function $S$ and the complementary sensitivity function $T$ satisfy their respective performance bounds over the entire bandwidth, confirming the feasibility of the robust specification.
@@ III.2 Simulation result
Following the completion of the controller synthesis, a high-fidelity Simulink-based DFACS simulation was developed for closed-loop validation, with the architecture and data flow shown in Figure 4.
@@
The simulation runs at 10 Hz over 25 days, initialized at 2035-01-01T00:00:00 UTC using DE440. Initial state perturbations of $\mathbf{p}$ are zero-mean Gaussian with $\sigma=5\times 10^{-8}\,\mathrm{m}$ for position variables and $5\times 10^{-8}\,\mathrm{rad}$ for attitude variables. The system incorporates actuation noises, environmental noises, and sensor readout noises, all based on the noise models published by [vidano2022lisa, NOISEchenglei2024review].
The DFACS simulation outputs both the true state $\mathbf{p}$ and its sensor readout $\tilde{\mathbf{p}}$. To assess compliance with the requirements in Table 1, the amplitude spectral densities (ASDs) of $\mathbf{r}_{m_{1}O_{1}}^{O_{1}}$, $\mathbf{r}_{m_{2}O_{2}}^{O_{2}}$, and $\mathbf{q}_{SC}$ are plotted in Figure 5.
@@
The above simulation results demonstrate that, within the measurement frequency band of 0.1 mHz to 1 Hz, the DFACS achieves the control requirements for each degree of freedom of $\mathbf{p}$ as specified in Table 1. Compared to the latest reported control performance of LISA [vidano2024ieee], these results show superior performance in the low-frequency band, which is attributed to the feedforward control design for the constant component of SRP.
@@ IV.1 Principle of CoM Offset Calibration in Science Mode
induced by the SC attitude jitter. Based on the simulation results, the amplitude spectral density of the acceleration data $\mathbf{\ddot{r}}_{m_{1}O_{1}}^{O_{1}}$ and $\mathbf{a}_{\dot{\boldsymbol{\omega}},1}$ is shown in Figures 6.
Figure 6: The amplitude spectral density plot of $\mathbf{\ddot{r}}_{m_{1}O_{1}}^{O_{1}}$ and $\mathbf{a}_{\dot{\omega},1}$, from left to right, corresponds to the x, y, and z-axis data, with the red line representing $\mathbf{\ddot{r}}_{m_{1}O_{1}}^{O_{1}}$ and the blue line representing $\mathbf{a}_{\dot{\omega},1}$.
@@
In contrast, during science mode calibration (i.e., without dedicated maneuvers), $\mathbf{a}_{\dot{\boldsymbol{\omega}},i}$ is neither dominant nor spectrally isolated, as evidenced by the broadband spectra in Figs. 6. Instead, Substituting Eq. (3) into Eq. (4) yields:
@@
In Eq. (36), $\mathbf{n}_{i}$ denotes the state vector of the shaped readout noise process. Since the Taiji mission has not yet released a comprehensive, publicly available characterization of its sensor noise transfer functions, we adopt the experimentally validated noise models from the LISA mission [vidano2022lisa]. This choice is justified by the high similarity between Taiji and LISA in scientific objectives, orbital environment, and core sensor technologies. Based on these established transfer-function models for the IS, IFO, and DWS, the corresponding state-space realizations for $\mathbf{n}_{i}$ can be constructed as
@@ IV.2 Addaptive Kalman filter
In previous work for maneuver-based COM offset calibration, algorithms such as least squares estimation or Kalman filtering [KF_ORI] have typically been used to estimate the CoM offset [wang2003study, COM_CALI_ZHANG, COM_CALI_WEI]. This paper continues this approach by using Kalman filtering for calibration. Based on Eq. (36), the following Kalman filter is constructed:
@@
where $\gamma$ represents the model parameter ratio of $\mathbf{Q}$ and $\mathbf{R}$. Given estimates $\hat{\mathbf{Q}}$ and $\hat{\mathbf{R}}$, the filter is optimal if and only if $\hat{\gamma}=\gamma$. For the present extended state model, $\gamma$ is 493-dimensional, making direct iterative tuning of $\hat{\mathbf{Q}}$ and $\hat{\mathbf{R}}$ impractical. To address this, we employ an adaptive Kalman filter based on the Sage-Husa method [akf_sage1969algorithms, akf_akhlaghi2017adaptive], which falls within the class of covariance matching techniques, a well-established approach for the estimation of process and measurement noise covariances [1972AKF_SAGEHUSA, 1976AKF_SAGEHUSA].
@@
where $\alpha\in(0,1)$ is the adaptation gain. The updating procedure in Eqs. (44)–(45) constitutes the innovation-based covariance matching method. However, this approach does not guarantee the positive definiteness of $\hat{\mathbf{Q}}_{k}$ and $\hat{\mathbf{R}}_{k}$, which may lead to filter divergence in high-dimensional settings. To enforce positive definiteness, the residual-based covariance matching method is often adopted [song2020AKF_COMATCHING]:
@@
To resolve this, we retain the monotonic update for $\hat{\mathbf{Q}}_{k}$ (critical for process-noise stability) but relax it for $\hat{\mathbf{R}}_{k}$, exploiting the availability of reliable a priori sensor noise characterization. The forgetting factor $(1-\alpha)$ is thus shifted to the innovation residual computation:
@@
$\hat{\mathbf{R}}_{k}=\alpha\,\hat{\mathbf{R}}_{k-1}+\min\left\{|\hat{\mathbf{R}}_{\mathrm{iter}}|,\,\hat{\mathbf{R}}_{k-1}\right\}$ (54)
@@
Figure 7 summarizes the CoM offset calibration procedure by ES-AKF based on Eqs.(55)-(63). For comparison, its performance is evaluated against the innovation-based and residual-based methods on the TM1–SC relative displacement readout $\tilde{\mathbf{r}}_{m_{1}O_{1}}^{O_{1}}$. Figures 1010 compare the filtering performance of three adaptive Kalman filter variants: the innovation-based method, the residual-based method, and the proposed approach. These are evaluated on the TM1–SC relative displacement $\mathbf{r}_{m_{1}O_{1}}^{O_{1}}$ along its three axes. The plots show the true trajectory (red), the raw readout data (yellow), the innovation-based estimate (blue), the residual-based estimate (green), and the proposed estimate (black).

### v2

@@ III.1 Controller design
The $H_{\infty}$ control method is a robust synthesis approach rooted in norm-optimal control theory, specifically formulated to ensure closed-loop stability and prescribed performance bounds in the presence of model uncertainties and exogenous disturbances. As illustrated in Fig. 2, the generalized plant $P$ represents the controlled dynamics (e.g., SC–TM coupled system), while $K$ denotes the to-be-designed controller. The objective is to synthesize a stabilizing controller $K$ that minimizes the worst-case gain from exogenous inputs $\mathbf{w}$ (e.g., disturbances, sensor noises) to regulated outputs $\mathbf{z}$ (e.g., TM–SC relative motion, attitude error). This is formalized as the $H_{\infty}$ norm minimization problem [20]
@@
Since analytical $H_{\infty}$ solutions are generally intractable, a suboptimal controller is obtained by minimizing $\eta$ subject to
@@
which reduces to a convex linear matrix inequality feasibility problem via the Bounded Real Lemma [8].
The suboptimal $H_{\infty}$ problem in Eq. (16) can be solved by Mixed Sensitivity Design [50]. Its inputs are the plant $G$ and weighting functions $W_{1}$, $W_{2}$, $W_{3}$, representing performance, control effort, and robustness requirements, respectively. The synthesis ensures the frequency-dependent bounds
@@
Where the sensitivity and complementary sensitivity functions are defined as $S=(I+GK)^{-1}$ and $T=GK(I+GK)^{-1}$. Enforcing an upper bound on $\|S\|$ improves low-frequency disturbance rejection and reference tracking; bounding $\|T\|$ ensures high-frequency robustness against sensor noise and unmodeled dynamics; and limiting $\|KS\|$ constrains control effort.Since $H_{\infty}$ synthesis via Mixed Sensitivity Design requires a state-space realization, the nonlinear DFACS dynamics, i.e, Eq. (4), Eq. (7), and Eq. (11) are linearized about the nominal drag-free trajectory. The resulting linear time-invariant model is
@@
Where $\mathbf{x}=[\mathbf{p}^{\top}\ \mathbf{\dot{p}}^{\top}]^{\top}$ is the state vector comprising relative position $\mathbf{p}$ and velocity $\mathbf{\dot{p}}$. The state variables of position are defined as
@@
where $\mathbf{F}_{E_{i}}\in\mathbb{R}^{2}$ denotes the 2-D electrostatic force on the $i$-th TM in the plane orthogonal to its sensitive axis. $A_{k}$ and $A_{v}$ in Eq.(18) is given by
$A_{k}=\begin{bmatrix}I/2&0_{3\times 3}&0_{3\times 3}&0_{3\times 3}&0_{3\times 3}\\ 0_{3\times 3}&I&0_{3\times 3}&0_{3\times 3}&0_{3\times 3}\\ 0_{3\times 3}&0_{3\times 3}&I/2&0_{3\times 3}&0_{3\times 3}\\ 0_{3\times 3}&0_{3\times 3}&0_{3\times 3}&I&0_{3\times 3}\\ 0_{3\times 3}&0_{3\times 3}&0_{3\times 3}&0_{3\times 3}&I/2\\ \end{bmatrix}\in\mathbb{R}^{15\times 15},$
$A_{v}=\begin{bmatrix}0_{3\times 3}&0_{3\times 3}&0_{3\times 3}&0_{3\times 3}&0_{3\times 3}\\ 0_{3\times 3}&S_{TT}&S_{TR}&0_{3\times 3}&0_{3\times 3}\\ 0_{3\times 3}&S_{RT}&S_{RR}&0_{3\times 3}&0_{3\times 3}\\ 0_{3\times 3}&0_{3\times 3}&0_{3\times 3}&S_{TT}&S_{TR}\\ 0_{3\times 3}&0_{3\times 3}&0_{3\times 3}&S_{RT}&S_{RR}\\ \end{bmatrix}\in\mathbb{R}^{15\times 15}.$
Where $\mathbf{S}_{TT}$, $\mathbf{S}_{TR}$, $\mathbf{S}_{RT}$ and $\mathbf{S}_{RR}$ are the stiffness matrices [14].
@@
where $\mathbf{B}_{v}^{\dagger}\in\mathbb{R}^{16\times 15}$ denotes the pseudoinverse of the input matrix $\mathbf{B}_{v}\in\mathbb{R}^{15\times 16}$. Substituting Eq.(21) into the velocity dynamics $\ddot{\mathbf{p}}=\mathbf{A}_{v}\mathbf{p}+\mathbf{B}_{v}\mathbf{u}$ yields the fully decoupled closed-loop form
@@
To satisfy the drag-free control requirements, MPS is commanded to allocate its thrust in the SRF as follows: the $x$-axis thrust component generates the control force associated with the symmetric mode $x_{m_{1}O_{1}}^{O_{1}}+x_{m_{2}O_{2}}^{O_{2}}$, while the $y$-axis thrust component drives the antisymmetric mode $x_{m_{1}O_{1}}^{O_{1}}-x_{m_{2}O_{2}}^{O_{2}}$. Additionally, the $z$-axis thrust provides the control force corresponding to the transverse symmetric mode $z_{m_{1}O_{1}}^{O_{1}}+z_{m_{2}O_{2}}^{O_{2}}$.:
$\mathbf{F}_{T,x}$ $=B_{v,{1,4}}^{\dagger}\bigl(u_{d,4}+u_{d,10}\bigr)=B_{v,1,10}^{\dagger}\bigl(u_{d,4}+u_{d,10}\bigr).$ (24)
$\mathbf{F}_{T,y}$ $=B_{v,2,4}^{\dagger}\bigl(u_{d,4}-u_{d,10}\bigr)=B_{v,2,10}^{\dagger}\bigl(-u_{d,4}+u_{d,10}\bigr).$ (25)
$\mathbf{F}_{T,z}$ $=B_{v,3,6}^{\dagger}\bigl(u_{d,6}+u_{d,12}\bigr)=B_{v,3,12}^{\dagger}\bigl(u_{d,6}+u_{d,12}\bigr).$ (26)
Clearly, this allocation strategy imposes constraints on Eq. (23), analogous constraints for attitude channels are based on the SC and TM inertia tensors (e.g., torque allocation proportional to $\mathbf{J}_{S}^{-1}$, $\mathbf{J}_{m}^{-1}$). This convex quadratic program is solved using active-set algorithm.
After completing the decoupling, the design of the $H_{\infty}$ controller continues. Eq. (21) and Eq. (22) enable us to shift from designing a controller for $\mathbf{u}$ to that for the decoupled SISO system $\mathbf{u}_{d}$. For each degree of freedom in $\mathbf{p}$, only the associated control requirements and noise characteristics pertinent to that specific degree of freedom need to be considered in determining the weighting functions in Eq. (17). The detailed DFACS control requirements are given in Table 1.
@@
State variable Requirement
@@
$y_{m_{i}O_{i}}^{O_{i}},\,z_{m_{i}O_{i}}^{O_{i}}$ $1\times 10^{-8}\,\mathrm{m}/\sqrt{\mathrm{Hz}}$
$\varphi_{\mathrm{SC}},\,\theta_{\mathrm{SC}},\,\psi_{\mathrm{SC}}$ $1\times 10^{-8}\,\mathrm{rad}/\sqrt{\mathrm{Hz}}\times f_{s}$
$\varphi_{m_{i}O_{i}}$ $2\times 10^{-7}\,\mathrm{rad}/\sqrt{\mathrm{Hz}}\times f_{r}$
$\theta_{m_{i}O_{i}},\,\psi_{m_{i}O_{i}}$ $1\times 10^{-8}\,\mathrm{rad}/\sqrt{\mathrm{Hz}}\times f_{s}$
@@
Where $F_{u}$ is the transfer function determined by the exogenous disturbance model, $F_{y}$ is the transfer function determined by the measurement noise model, and $B$ is determined by the control requirements of each state variable. In the disturbance models, special attention is given to solar radiation pressure (SRP), with its transfer function given by [53]
@@
The constant $k_{d\mathrm{SC}}$ is derived from the SC’s orbital distance and the effective area of its solar arrays [19]. Given the similarity in orbital configuration ($\mathbf{r}_{SI}^{I}\approx 1~\text{AU}$) and spacecraft structural layout, we adopt the value used in the LISA mission studies, namely, $k_{d\mathrm{SC}}=6.3513\times 10^{-5}~\text{N},$ which is significantly larger than the magnitude of the transfer function $H_{\mathrm{press}}$ ($\sim 1\times 10^{-10}~\text{N}$). Therefore, including SRP directly in $F_{u}$ would significantly increase the requirements on $W_{1}$ in the low-frequency range, making Eq. (17) difficult to solve. To address this issue, we introduce a feedforward channel that leverages spacecraft orbit determination data to estimate $\hat{\mathbf{r}}_{SI}^{I}$. Based on Taiji’s exceptional orbital stability and its periodic ground-based orbit determination architecture, a conservative 50-km $1\sigma$ position estimation error is assumed for $\hat{\mathbf{r}}_{SI}^{I}$ [26]. This allows for the estimation of the ultra-low-frequency components of $\mathbf{F}_{SRP}$ and $\mathbf{D}_{SRP}$, denoted as $\hat{\mathbf{F}}_{SRP,C}$ and $\hat{\mathbf{D}}_{SRP,C}$
@@
Fig. 3 shows the mixed sensitivity $H_{\infty}$ controller design results. Critically, both the sensitivity function $S$ and the complementary sensitivity function $T$ satisfy their respective performance bounds over the entire bandwidth, confirming the feasibility of the robust specification.
@@ III.2 Simulation result
Following the completion of the controller synthesis, a high-fidelity DFACS simulation was developed for closed-loop validation, with the architecture and data flow shown in Fig. 4.
@@
The simulation runs at 10 Hz over 25 days, initialized at 2035-01-01T00:00:00 UTC using the planetary and lunar ephemerides DE440. Initial state perturbations of $\mathbf{p}$ are zero-mean Gaussian with $\sigma=5\times 10^{-8}\,\mathrm{m}$ for position variables and $5\times 10^{-8}\,\mathrm{rad}$ for attitude variables.
The DFACS simulation outputs both the true state $\mathbf{p}$ and its sensor readout $\tilde{\mathbf{p}}$. To assess compliance with the requirements in Table 1, the amplitude spectral densities (ASDs) of $\mathbf{r}_{m_{1}O_{1}}^{O_{1}}$, $\mathbf{r}_{m_{2}O_{2}}^{O_{2}}$, and $\mathbf{q}_{SC}$ are plotted in Fig. 5.
Table 2: The Parameters of the simulation system
Parts Parameters Values
Spacecraft Mass 1500 kg
Inertias $J_{S}=\begin{bmatrix}800&13&10\\ 13&800&12\\ 10&12&1000\end{bmatrix}\ \mathrm{kg\cdot m^{2}}$
TM Mass 1.96 kg
Inertias $J_{m}=\begin{bmatrix}6.913&0&0\\ 0&6.913&0\\ 0&0&6.913\end{bmatrix}\times 10^{-4}\ \mathrm{kg\cdot m^{2}}$
SC-TM CoM offset $\mathbf{b}^{S}_{1}=[0.001,\ 0.18,\ 0]\ \mathrm{m}$ $\mathbf{b}^{S}_{2}=[0.001,\ -0.18,\ 0]\ \mathrm{m}$
/ Simulation time 25 days
@@
The above simulation results demonstrate that, within the measurement frequency band of 0.1 mHz to 1 Hz, the DFACS achieves the control requirements for each degree of freedom of $\mathbf{p}$ as specified in Table 1.
@@ IV.1 Principle of CoM Offset Calibration in Science Mode
induced by the SC attitude jitter. Based on the simulation results, the amplitude spectral density of the acceleration data $\mathbf{\ddot{r}}_{m_{1}O_{1}}^{O_{1}}$ and $\mathbf{a}_{\dot{\boldsymbol{\omega}},1}$ is shown in Figs. 6.
Figure 6: The amplitude spectral density plot of $\mathbf{\ddot{r}}_{m_{1}O_{1}}^{O_{1}}$ and $\mathbf{a}_{\dot{\omega},1}$, from top to bottom, corresponds to the x, y, and z-axis data, with the red line representing $\mathbf{\ddot{r}}_{m_{1}O_{1}}^{O_{1}}$ and the blue line representing $\mathbf{a}_{\dot{\omega},1}$.
@@
In contrast, during science mode calibration (i.e., without dedicated maneuvers), $\mathbf{a}_{\dot{\boldsymbol{\omega}},i}$ is neither dominant nor spectrally isolated, as evidenced by the broadband spectra in Fig. 6. Instead, substituting Eq. (3) into Eq. (4) yields:
@@
In Eq. (37), $\mathbf{n}_{i}$ denotes the state vector of the shaped readout noise. We adopt the experimentally validated noise models from the LISA mission [51]. Since the high similarity between Taiji and LISA in scientific objectives, orbital environment, and payload configurations. Based on these established transfer-function models for the IS, IFO, and DWS, the corresponding state-space realizations for $\mathbf{n}_{i}$ can be constructed as
@@ IV.2 Addaptive Kalman filter
In previous work for maneuver-based CoM offset calibration, algorithms such as least squares estimation or Kalman filtering [23] have been used [55, 63, 58]. Based on Eq. (37), the following Kalman filter is constructed:
@@
where $\gamma$ represents the model parameter ratio of $\mathbf{Q}$ and $\mathbf{R}$. Given estimates $\hat{\mathbf{Q}}$ and $\hat{\mathbf{R}}$, the filter is optimal if and only if $\hat{\gamma}=\gamma$. For the present extended state model, $\gamma$ is 493-dimensional, making direct iterative tuning of $\hat{\mathbf{Q}}$ and $\hat{\mathbf{R}}$ impractical. To address this, we employ an adaptive Kalman filter based on the Sage-Husa method [43, 3], which falls within the class of covariance matching techniques, a well-established approach for the estimation of process and measurement noise covariances [34].
@@
where $\alpha\in(0,1)$ is the adaptation gain. The updating procedure in Eqs. (45)–(46) constitutes the innovation-based covariance matching method. However, this approach does not guarantee the positive definiteness of $\hat{\mathbf{Q}}_{k}$ and $\hat{\mathbf{R}}_{k}$, which may lead to filter divergence in high-dimensional cases. To enforce positive definiteness, the residual-based covariance matching method is often adopted [47]:
@@
To resolve this, we retain the monotonic update for $\hat{\mathbf{Q}}_{k}$ (critical for process-noise stability) but relax it for $\hat{\mathbf{R}}_{k}$, exploiting the availability of reliable a priori sensor noise characterization. The factor $(1-\alpha)$ is thus shifted to the innovation residual computation,
@@
$\hat{\mathbf{R}}_{k}=\alpha\,\hat{\mathbf{R}}_{k-1}+\min\left\{|\hat{\mathbf{R}}_{\mathrm{iter}}|,\,\hat{\mathbf{R}}_{k-1}\right\}.$ (55)
@@
Based on Eqs.(56)-(64), the extended state $\mathbf{X}_{i}$ is estimated using the proposed ES-AKF. For comparison, its performance is evaluated against the innovation-based and residual-based methods on the TM1–SC relative displacement readout $\tilde{\mathbf{r}}_{m_{1}O_{1}}^{O_{1}}$. Figs. 99 compare the filtering performance of three adaptive Kalman filter variants: the innovation-based method, the residual-based method, and the proposed approach. These are evaluated on the TM1–SC relative displacement $\mathbf{r}_{m_{1}O_{1}}^{O_{1}}$ along its three measurement axes. The plots show the true trajectory (red), the raw readout data (yellow), the innovation-based estimate (blue), the residual-based estimate (green), and the proposed estimate (black).